OpenStudy (ayabaya2002):

Variable y varies directly with variable x, and y = 5 when x = 8. What is the value of x when y = 15?

OpenStudy (cutiekawdfi):

8.3

OpenStudy (ayabaya2002):

ur good at this

OpenStudy (faiqraees):

the answer is 24

OpenStudy (faiqraees):

y varies directly with x y=kx 5=k8 k=5/8 y=5/8 x 15=5/8 x x=24

OpenStudy (faiqraees):

or more simply when y = 5 x=8 y varies directly with x so if y is multiplied by 3 so is x y=15 x=24